# Fast relaxation of a viscous vortex in an external flow

### v1

@@ 4.2.1 Relaxation for small time
$\|\tilde{w}(\tau)\|_{\mathcal{Y}}\,\leq\,C_{4}\Bigl{(}\frac{\varepsilon^{5}}{\delta^{2}}+\varepsilon^{2}\Bigr{)}+C_{4}\,e^{\tau}\,\delta^{1/6}\Bigl{(}\log\frac{2}{\delta}\Bigr{)}^{1/2}\Bigl{(}\frac{t_{0}}{T_{0}}\Bigr{)}^{2}\,,$ (4.19)
for all $\tau\in\bigl{[}0,\log(T/t_{0})\bigr{]}$.
@@
What remains is to estimate all the terms in (4.20) that involve the solution $w_{0}$ of the linear equation (4.5). We start with the advection term $\mathcal{A}[w_{0}]$. Integrating by parts, we obtain
@@
$\|v_{0}\|_{L^{\infty}}\,\leq\,C\|w_{0}\|_{L^{1}\cap L^{2}}\Bigl{(}1+\log_{+}\frac{\|w_{0}\|_{L^{3}}}{\|w_{0}\|_{L^{1}\cap L^{2}}}\Bigr{)}^{1/2}\,\leq\,C\|w_{0}\|_{\mathcal{Y}}\Bigl{(}\log\frac{2}{\delta}\Bigr{)}^{1/2}\,,$
where in the last inequality we used the a priori estimate on $\|w_{0}\|_{L^{3}}$ given by Lemma 4.4. We arrive at
$\bigl{|}\langle\tilde{w},\mathcal{B}[w_{0},w_{0}+\tilde{w}]\rangle_{\mathcal{Y}}\bigr{|}\,\leq\,C\Bigl{(}\log\frac{2}{\delta}\Bigr{)}^{1/2}\mathcal{F}[\tilde{w}]^{1/2}\|w_{0}\|_{\mathcal{Y}}\bigl{(}\|w_{0}\|_{\mathcal{Y}}+\|\tilde{w}\|_{\mathcal{Y}}\bigr{)}\,.$ (4.23)
The second case to consider is $w_{1}=\tilde{w}$ and $w_{2}=w_{0}$. Here we invoke [5, Lemma 5.6] which directly gives
$\|\tilde{v}w_{0}\|_{\mathcal{Y}}\,\leq\,C\|w_{0}\|_{\mathcal{Y}}\|\tilde{w}\|_{L^{1}\cap L^{2}}\Bigl{(}1+\log_{+}\frac{\|p^{1/2}w_{0}\|_{L^{3}}}{\|w_{0}\|_{\mathcal{Y}}}\Bigr{)}^{1/2}\,\leq\,C\|w_{0}\|_{\mathcal{Y}}\|\tilde{w}\|_{\mathcal{Y}}\Bigl{(}\log\frac{2}{\delta}\Bigr{)}^{1/2}\,,$
@@
$\bigl{|}\langle\tilde{w},\mathcal{B}[\tilde{w},w_{0}]\rangle_{\mathcal{Y}}\bigr{|}\,\leq\,C\Bigl{(}\log\frac{2}{\delta}\Bigr{)}^{1/2}\mathcal{F}[\tilde{w}]^{1/2}\|w_{0}\|_{\mathcal{Y}}\|\tilde{w}\|_{\mathcal{Y}}\,.$ (4.24)
@@
$+\,\frac{Ct}{T_{0}}\,\mathcal{F}[\tilde{w}]^{1/2}\bigl{(}\|w_{0}\|_{\mathcal{Y}}+\||\xi|w_{0}\|_{\mathcal{Y}}\bigr{)}+C\Bigl{(}\log\frac{2}{\delta}\Bigr{)}^{1/2}\mathcal{F}[\tilde{w}]^{1/2}\|w_{0}\|_{\mathcal{Y}}\bigl{(}\|w_{0}\|_{\mathcal{Y}}+\|\tilde{w}\|_{\mathcal{Y}}\bigr{)}\,,$
@@
$\mathcal{M}(\tau)\,$ $=\,\frac{Ct}{T_{0}}+C\Bigl{(}\log\frac{2}{\delta}\Bigr{)}\|w_{0}(\tau)\|_{\mathcal{Y}}^{2}\,,$
$\mathcal{S}(\tau)\,$ $=\,C\Bigl{(}\frac{\varepsilon^{5}}{\delta^{2}}+\varepsilon^{2}\Bigr{)}^{2}+C\Bigl{(}\frac{t}{T_{0}}\Bigr{)}^{2}\Bigl{(}\|w_{0}(\tau)\|_{\mathcal{Y}}^{2}+\||\xi|w_{0}(\tau)\|_{\mathcal{Y}}^{2}\Bigr{)}+C\Bigl{(}\log\frac{2}{\delta}\Bigr{)}\|w_{0}(\tau)\|_{\mathcal{Y}}^{4}\,.$
@@
$\int_{0}^{\log(T/t_{0})}\mathcal{M}(\tau)\,\mathrm{d}\tau\,\leq\,\frac{CT}{T_{0}}+C\delta^{1/3}\Bigl{(}\log\frac{2}{\delta}\Bigr{)}\|\phi_{0}\|_{\mathcal{Y}}^{2}\,\leq\,K\,,$
@@
$\int_{0}^{\tau}\mathcal{S}(\tau^{\prime})\,\mathrm{d}\tau^{\prime}\,$ $\leq\,C\Bigl{(}\frac{\varepsilon^{5}}{\delta^{2}}+\varepsilon^{2}\Bigr{)}^{2}+C\Bigl{(}\frac{t}{T_{0}}\Bigr{)}^{2}\,\delta^{1/3}\Bigl{(}\log\frac{2}{\delta}\Bigr{)}\|\phi_{0}\|_{\gamma}^{2}+C\,\delta^{1/3}\Bigl{(}\log\frac{2}{\delta}\Bigr{)}\|\phi_{0}\|_{\mathcal{Y}}^{4}$
$\leq\,C\Bigl{(}\frac{\varepsilon^{5}}{\delta^{2}}+\varepsilon^{2}\Bigr{)}^{2}+C\,\delta^{1/3}\Bigl{(}\log\frac{2}{\delta}\Bigr{)}\Bigl{(}\frac{t\,t_{0}}{T_{0}^{2}}\Bigr{)}^{2}\,,$
because $\|\phi_{0}\|_{\mathcal{Y}}\leq C\|\phi_{0}\|_{\gamma}\leq Ct_{0}/T_{0}\leq Ct/T_{0}$. So, applying Grönwall’s lemma to the differential inequality (4.25), we obtain
$\mathcal{E}[\tilde{w}(\tau)]\,\leq\,e^{K}\int_{0}^{\tau}\mathcal{S}(\tau^{\prime})\,\mathrm{d}\tau^{\prime}\,\leq\,C\Bigl{(}\frac{\varepsilon^{5}}{\delta^{2}}+\varepsilon^{2}\Bigr{)}^{2}+C\,\delta^{1/3}\Bigl{(}\log\frac{2}{\delta}\Bigr{)}\Bigl{(}\frac{t\,t_{0}}{T_{0}^{2}}\Bigr{)}^{2}\,.$ (4.26)
If $\delta$ is sufficiently small, this ensures that the a priori estimate $C_{0}\mathcal{E}[\tilde{w}]^{1/2}\leq\kappa/2$ holds whenever $\tau\leq\log(T/t_{0})$. Finally, since $\|\tilde{w}(\tau)\|_{\mathcal{Y}}=\mathcal{E}[\tilde{w}(\tau)]^{1/2}$ and $t=t_{0}e^{\tau}$, we see that (4.19) follows from (4.26). ∎
@@ 4.2.2 Relaxation for large time
^$\hat{\mathcal{E}}(\tau)^{1/2}\,\leq\,C_{5}\Bigl{(}\frac{\varepsilon^{5}}{\delta^{2}}+\varepsilon^{2}\Bigr{)}+C_{5}\,e^{\tau}\,\delta^{1/6}\Bigl{(}\log\frac{2}{\delta}\Bigr{)}^{1/2}\Bigl{(}\frac{t_{0}}{T_{0}}\Bigr{)}^{2}\,,$ (4.27)
for all $\tau\in\bigl{[}0,\log(T/t_{0})\bigr{]}$.
@@
$|\mathcal{G}(\tau)|\,\leq\,C\hat{\mathcal{F}}(\tau)^{1/2}\Bigl{\{}\|w_{0}\|_{\mathcal{Y}}+\||\xi|w_{0}\|_{\mathcal{Y}}+\Bigl{(}\log\frac{2}{\delta}\Bigr{)}^{1/2}\|w_{0}\|_{\mathcal{Y}}\bigl{(}\|w_{0}\|_{\mathcal{Y}}+\|\tilde{w}\|_{\mathcal{Y}}\bigr{)}\Bigr{\}}\,.$ (4.29)
@@
$\delta\,\|\tilde{w}(\cdot,\tau)\|_{L^{1}}\,\leq\,C\delta\,\|\tilde{w}(\cdot,\tau)\|_{\mathcal{Y}}\,=\,C\delta\,\hat{\mathcal{E}}(\tau)^{1/2}\,\leq\,C\varepsilon^{2}\bigl{(}\varepsilon+\delta\bigr{)}+C\varepsilon^{2}\,\delta^{1/6}\Bigl{(}\log\frac{2}{\delta}\Bigr{)}^{1/2}\frac{t_{0}}{T_{0}}\,,$
where we used again the relation $\varepsilon^{2}=\delta t/T_{0}$. Altogether arrive at
$\mathcal{I}_{1}(t)\,\leq\,C\varepsilon^{2}\biggl{\{}\varepsilon+\delta+\delta^{1/6}\Bigl{(}\log\frac{2}{\delta}\Bigr{)}^{1/2}\frac{t_{0}}{T_{0}}+\Bigl{(}\frac{t_{0}}{t}\Bigr{)}^{\beta+1}\biggr{\}}\,,\qquad t\in(t_{0},T)\,,$ (4.30)
and the second integral that appears in the the proof of Theorem 1.6 satisfies $\mathcal{I}_{2}(t)\leq C\varepsilon^{2}(\varepsilon+\delta)$, so that $\mathcal{I}_{2}(t)\leq\mathcal{I}_{1}(t)$. Estimate (1.17) thus follows from (4.30). ∎
@@ A.1 Comparison with the Burgers vortex
This means that, for $t\in(0,T_{0})$ and $\delta>0$ sufficiently small, the approximate solution (1.10) is a simple rescaling of the vorticity profile of a Burgers vortex with asymmetry parameter $\gamma t\in(0,1)$ and Reynolds number $1/\delta\gg 1$.
@@ A.2 Partial inverse for the advection operator $\Lambda$
In this section, for completeness, we recall the known formulas for the (partial) inverse of the integro-differential operator $\Lambda$ defined in (2.14). More details can be found in the references [4, 6]. Since $\Lambda$ leaves invariant the direct sum decomposition (2.12), it is sufficient to study the restriction of $\Lambda$ to each subspace $\mathcal{Y}_{n}$. Actually $\mathcal{Y}_{0}\subset\mathrm{Ker}(\Lambda)$ by (2.15), so we can assume that $n\geq 1$. To exploit the rotational symmetry, we use polar coordinates in $\mathbb{R}^{2}$ defined by $\xi=(r\cos\theta,r\sin\theta)$, and we introduce the radially symmetric functions
$v_{0}(r)\,=\,\frac{1}{2\pi r^{2}}\bigl{(}1-e^{-r^{2}/4}\bigr{)}\,,\qquad g(r)\,=\,\frac{1}{8\pi}\,e^{-r^{2}/4}\,,\qquad h(r)\,=\,\frac{g(r)}{v_{0}(r)}\,=\,\frac{r^{2}/4}{e^{r^{2}/4}-1}\,.$ (A.5)
Note that $U_{0}=v\xi^{\perp}$ and $\nabla\Omega_{0}=-g\xi$, where $\Omega_{0}$ and $U_{0}$ are defined in (1.6).
@@
$\Lambda\Omega\,=\,U_{0}\cdot\nabla\Omega+U\cdot\nabla\Omega_{0}\,=\,n\bigl{(}vw-g\varphi\bigr{)}\sin(n\theta)\,.$ (A.8)
Similarly, if $\Omega=w(r)\sin(n\theta)$, then $\Lambda\Omega=n\bigl{(}vw-g\varphi\bigr{)}\cos(n\theta)$.
@@
$-\varphi^{\prime\prime}(r)-\frac{1}{r}\,\varphi^{\prime}(r)+\Bigl{(}\frac{n^{2}}{r^{2}}-h(r)\Bigr{)}\varphi(r)\,=\,\frac{b(r)}{nv_{0}(r)}\,,\qquad r>0\,,$ (A.9)
@@
$w(r)\,=\,\varphi(r)h(r)+\frac{b(r)}{nv_{0}(r)}\,,\qquad r>0\,,$ (A.10)
which is depicted in Figure 2 in the case $n=2$. Then (A.7) is obviously satisfied, and (A.8) implies that $\Lambda\Omega=F$. The same conclusion holds if $F=b(r)\cos(n\theta)$ and $\Omega=w(r)\sin(n\theta)$. The level lines of $\Omega$ in the case $n=2$ are depicted on the right of Figure 3.
@@
If $n=1$, the homogeneous differential equation (A.9) with $b=0$ has a nontrivial solution $\varphi=rv$ which satisfies the boundary conditions. As a consequence, the inhomogeneous equation can be solved only if the right-hand side satisfies $\int_{0}^{\infty}b(r)r^{2}\,\mathrm{d}r=0$, and the solution is never unique. The solvability condition ensures that $F$ belongs to the subspace $\mathcal{Y}_{1}^{\prime}$ defined by

### v2

@@ 4.2.1 Relaxation for small time
$\|\tilde{w}(\tau)\|_{\mathcal{Y}}\,\leq\,C_{4}\Bigl{(}\frac{\varepsilon^{5}}{\delta^{2}}+\varepsilon^{2}\Bigr{)}+C_{4}\,\delta^{1/6}\Bigl{(}\log\frac{2}{\delta}\Bigr{)}^{1/2}\,\frac{t}{T_{0}}\,\Theta\Bigl{(}\frac{t_{0}}{T_{0}}\Bigr{)}\,,$ (4.19)
for all $\tau\in\bigl{[}0,\log(T/t_{0})\bigr{]}$, where $\Theta(s)=s\bigl{(}1+\log_{+}(1/s)\bigr{)}^{1/2}$ and $\log_{+}(s)=\max(\log(s),0)$.
@@
It remains to estimate all the terms in (4.20) that involve the solution $w_{0}$ of the linear equation (4.5). We start with the advection term $\mathcal{A}[w_{0}]$. Integrating by parts, we obtain
@@
$\|v_{0}\|_{L^{\infty}}\,\leq\,C\|w_{0}\|_{L^{1}\cap L^{2}}\Bigl{(}1+\log_{+}\frac{\|w_{0}\|_{L^{3}}}{\|w_{0}\|_{L^{1}\cap L^{2}}}\Bigr{)}^{1/2}\,\leq\,C\Bigl{(}\log\frac{2}{\delta}\Bigr{)}^{1/2}\,\Theta\bigl{(}\|w_{0}\|_{\mathcal{Y}}\bigr{)}\,,$
where in the second inequality we used the fact that $\|w_{0}\|_{L^{1}\cap L^{2}}\leq C\|w_{0}\|_{\mathcal{Y}}$ and $\|w_{0}\|_{L^{3}}\leq C\delta^{-1}$, see (4.14). We deduce that
$\bigl{|}\langle\tilde{w},\mathcal{B}[w_{0},w_{0}+\tilde{w}]\rangle_{\mathcal{Y}}\bigr{|}\,\leq\,C\Bigl{(}\log\frac{2}{\delta}\Bigr{)}^{1/2}\mathcal{F}[\tilde{w}]^{1/2}\,\Theta\bigl{(}\|w_{0}\|_{\mathcal{Y}}\bigr{)}\bigl{(}\|w_{0}\|_{\mathcal{Y}}+\|\tilde{w}\|_{\mathcal{Y}}\bigr{)}\,.$ (4.23)
The second case is $w_{1}=\tilde{w}$ and $w_{2}=w_{0}$. Here we invoke [5, Lemma 5.6] which gives
$\|\tilde{v}w_{0}\|_{\mathcal{Y}}\,\leq\,C\|w_{0}\|_{\mathcal{Y}}\|\tilde{w}\|_{L^{1}\cap L^{2}}\Bigl{(}1+\log_{+}\frac{\|p^{1/2}w_{0}\|_{L^{3}}}{\|w_{0}\|_{\mathcal{Y}}}\Bigr{)}^{1/2}\,\leq\,C\Bigl{(}\log\frac{2}{\delta}\Bigr{)}^{1/2}\Theta\bigl{(}\|w_{0}\|_{\mathcal{Y}}\bigr{)}\|\tilde{w}\|_{\mathcal{Y}}\,,$
@@
$\bigl{|}\langle\tilde{w},\mathcal{B}[\tilde{w},w_{0}]\rangle_{\mathcal{Y}}\bigr{|}\,\leq\,C\Bigl{(}\log\frac{2}{\delta}\Bigr{)}^{1/2}\mathcal{F}[\tilde{w}]^{1/2}\,\Theta\bigl{(}\|w_{0}\|_{\mathcal{Y}}\bigr{)}\|\tilde{w}\|_{\mathcal{Y}}\,.$ (4.24)
@@
$+\,\frac{Ct}{T_{0}}\,\mathcal{F}[\tilde{w}]^{1/2}\bigl{(}\|w_{0}\|_{\mathcal{Y}}+\||\xi|w_{0}\|_{\mathcal{Y}}\bigr{)}+C\Bigl{(}\log\frac{2}{\delta}\Bigr{)}^{1/2}\mathcal{F}[\tilde{w}]^{1/2}\Theta\bigl{(}\|w_{0}\|_{\mathcal{Y}}\bigr{)}\bigl{(}\|w_{0}\|_{\mathcal{Y}}+\|\tilde{w}\|_{\mathcal{Y}}\bigr{)}\,,$
@@
$\mathcal{M}(\tau)\,$ $=\,\frac{Ct}{T_{0}}+C\Bigl{(}\log\frac{2}{\delta}\Bigr{)}\Theta\bigl{(}\|w_{0}(\tau)\|_{\mathcal{Y}}\bigr{)}^{2}\,,$
$\mathcal{S}(\tau)\,$ $=\,C\Bigl{(}\frac{\varepsilon^{5}}{\delta^{2}}+\varepsilon^{2}\Bigr{)}^{2}\!+C\Bigl{(}\frac{t}{T_{0}}\Bigr{)}^{2}\Bigl{(}\|w_{0}(\tau)\|_{\mathcal{Y}}^{2}+\||\xi|w_{0}(\tau)\|_{\mathcal{Y}}^{2}\Bigr{)}$
$\quad\,+C\Bigl{(}\log\frac{2}{\delta}\Bigr{)}\Theta\bigl{(}\|w_{0}(\tau)\|_{\mathcal{Y}}\bigr{)}^{2}\|w_{0}(\tau)\|_{\mathcal{Y}}^{2}\,.$
@@
$\int_{0}^{\log(T/t_{0})}\mathcal{M}(\tau)\,\mathrm{d}\tau\,\leq\,\frac{CT}{T_{0}}+C\delta^{1/3}\Bigl{(}\log\frac{2}{\delta}\Bigr{)}\Theta\bigl{(}\|\phi_{0}\|_{\mathcal{Y}}\bigr{)}^{2}\,\leq\,K\,,$
@@
$\int_{0}^{\tau}\mathcal{S}(\tau^{\prime})\,\mathrm{d}\tau^{\prime}\,$ $\leq\,C\Bigl{(}\frac{\varepsilon^{5}}{\delta^{2}}+\varepsilon^{2}\Bigr{)}^{2}+C\Bigl{(}\frac{t}{T_{0}}\Bigr{)}^{2}\,\delta^{1/3}\Bigl{(}\log\frac{2}{\delta}\Bigr{)}\|\phi_{0}\|_{\gamma}^{2}+C\,\delta^{1/3}\Bigl{(}\log\frac{2}{\delta}\Bigr{)}\Theta\bigl{(}\|\phi_{0}\|_{\mathcal{Y}}\bigr{)}^{2}\|\phi_{0}\|_{\mathcal{Y}}^{2}$
$\leq\,C\Bigl{(}\frac{\varepsilon^{5}}{\delta^{2}}+\varepsilon^{2}\Bigr{)}^{2}+C\,\delta^{1/3}\Bigl{(}\log\frac{2}{\delta}\Bigr{)}\Bigl{(}\frac{t}{T_{0}}\Bigr{)}^{2}\Theta\Bigl{(}\frac{t_{0}}{T_{0}}\Bigr{)}^{2}\,,$
because $\|\phi_{0}\|_{\mathcal{Y}}\leq C\|\phi_{0}\|_{\gamma}\leq Ct_{0}/T_{0}$ and $t_{0}\leq t$. So, applying Grönwall’s lemma to the differential inequality (4.25), we obtain
$\mathcal{E}[\tilde{w}(\tau)]\,\leq\,e^{K}\int_{0}^{\tau}\mathcal{S}(\tau^{\prime})\,\mathrm{d}\tau^{\prime}\,\leq\,C\Bigl{(}\frac{\varepsilon^{5}}{\delta^{2}}+\varepsilon^{2}\Bigr{)}^{2}+C\,\delta^{1/3}\Bigl{(}\log\frac{2}{\delta}\Bigr{)}\Bigl{(}\frac{t}{T_{0}}\Bigr{)}^{2}\Theta\Bigl{(}\frac{t_{0}}{T_{0}}\Bigr{)}^{2}\,.$ (4.26)
If $\delta$ is sufficiently small, this ensures that the a priori estimate $C_{0}\mathcal{E}[\tilde{w}]^{1/2}\leq\kappa/2$ holds whenever $\tau\leq\log(T/t_{0})$. Finally, since $\|\tilde{w}(\tau)\|_{\mathcal{Y}}=\mathcal{E}[\tilde{w}(\tau)]^{1/2}$ we see that (4.19) follows from (4.26). ∎
@@ 4.2.2 Relaxation for large time
^$\hat{\mathcal{E}}(\tau)^{1/2}\,\leq\,C_{5}\Bigl{(}\frac{\varepsilon^{5}}{\delta^{2}}+\varepsilon^{2}\Bigr{)}+C_{5}\,\delta^{1/6}\Bigl{(}\log\frac{2}{\delta}\Bigr{)}^{1/2}\,\frac{t}{T_{0}}\,\Theta\Bigl{(}\frac{t_{0}}{T_{0}}\Bigr{)}\,,$ (4.27)
for all $\tau\in\bigl{[}0,\log(T/t_{0})\bigr{]}$, where $\Theta(s)=s\bigl{(}1+\log_{+}(1/s)\bigr{)}^{1/2}$.
@@
$|\mathcal{G}(\tau)|\,\leq\,C\hat{\mathcal{F}}(\tau)^{1/2}\biggl{\{}\frac{t}{T_{0}}\bigl{(}\|w_{0}\|_{\mathcal{Y}}+\||\xi|w_{0}\|_{\mathcal{Y}}\bigr{)}+\Bigl{(}\log\frac{2}{\delta}\Bigr{)}^{1/2}\Theta\bigl{(}\|w_{0}\|_{\mathcal{Y}}\bigr{)}\bigl{(}\|w_{0}\|_{\mathcal{Y}}+\|\tilde{w}\|_{\mathcal{Y}}\bigr{)}\biggr{\}}\,.$ (4.29)
@@
$\delta\,\|\tilde{w}(\cdot,\tau)\|_{L^{1}}\,\leq\,C\delta\,\|\tilde{w}(\cdot,\tau)\|_{\mathcal{Y}}\,=\,C\delta\,\hat{\mathcal{E}}(\tau)^{1/2}\,\leq\,C\varepsilon^{2}\bigl{(}\varepsilon+\delta\bigr{)}+C\varepsilon^{2}\,\delta^{1/6}\Bigl{(}\log\frac{2}{\delta}\Bigr{)}^{1/2}\Theta\Bigl{(}\frac{t_{0}}{T_{0}}\Bigr{)}\,,$
where we used again the relation $\varepsilon^{2}=\delta t/T_{0}$. We thus arrive at
$\mathcal{I}_{1}(t)\,\leq\,C\varepsilon^{2}\biggl{\{}\varepsilon+\delta+\delta^{1/6}\Bigl{(}\log\frac{2}{\delta}\Bigr{)}^{1/2}\Theta\Bigl{(}\frac{t_{0}}{T_{0}}\Bigr{)}+\Bigl{(}\frac{t_{0}}{t}\Bigr{)}^{\beta+1}\biggr{\}}\,,\qquad t\in(t_{0},T)\,,$
and the second integral that appears in the proof of Theorem 1.6 satisfies $\mathcal{I}_{2}(t)\leq C\varepsilon^{2}(\varepsilon+\delta)$. Altogether we obtain estimate (1.17). ∎
@@ A.1 Comparison with the Burgers vortex
This means that, for $t\in(0,T_{0})$ and $\delta>0$ sufficiently small, the approximate solution (1.10) is essentially a rescaling of the vorticity profile of a Burgers vortex with asymmetry parameter $\gamma t\in(0,1)$ and Reynolds number $1/\delta\gg 1$.
@@ A.2 Partial inverse for the advection operator $\Lambda$
In this section, for completeness, we recall the known formulas for the (partial) inverse of the integro-differential operator $\Lambda$ defined in (2.14). More details can be found in the references [4, 6]. Since $\Lambda$ leaves invariant the direct sum decomposition (2.12), it is sufficient to study the restriction of $\Lambda$ to each subspace $\mathcal{Y}_{n}$. Actually $\mathcal{Y}_{0}\subset\mathrm{Ker}(\Lambda)$ by (2.15), so we can assume that $n\geq 1$. To exploit the rotational symmetry, we use polar coordinates in $\mathbb{R}^{2}$ defined by $\xi=(r\cos\theta,r\sin\theta)$, and we consider the radially symmetric functions
$v_{*}(r)\,=\,\frac{1}{2\pi r^{2}}\bigl{(}1-e^{-r^{2}/4}\bigr{)}\,,\qquad g(r)\,=\,\frac{1}{8\pi}\,e^{-r^{2}/4}\,,\qquad h(r)\,=\,\frac{g(r)}{v_{*}(r)}\,=\,\frac{r^{2}/4}{e^{r^{2}/4}-1}\,.$ (A.5)
Note that $U_{0}(\xi)=v_{*}(|\xi|)\xi^{\perp}$ and $\nabla\Omega_{0}(\xi)=-g(|\xi|)\xi$, where $\Omega_{0}$ and $U_{0}$ are defined in (1.6).
@@
$\Lambda\Omega\,=\,U_{0}\cdot\nabla\Omega+U\cdot\nabla\Omega_{0}\,=\,n\bigl{(}v_{*}w-g\varphi\bigr{)}\sin(n\theta)\,.$ (A.8)
Similarly, if $\Omega=w(r)\sin(n\theta)$, then $\Lambda\Omega=n\bigl{(}v_{*}w-g\varphi\bigr{)}\cos(n\theta)$.
@@
$-\varphi^{\prime\prime}(r)-\frac{1}{r}\,\varphi^{\prime}(r)+\Bigl{(}\frac{n^{2}}{r^{2}}-h(r)\Bigr{)}\varphi(r)\,=\,\frac{b(r)}{nv_{*}(r)}\,,\qquad r>0\,,$ (A.9)
@@
$w(r)\,=\,\varphi(r)h(r)+\frac{b(r)}{nv_{*}(r)}\,,\qquad r>0\,,$ (A.10)
Then (A.7) is obviously satisfied, and (A.8) implies that $\Lambda\Omega=F$. The same conclusion holds if $F=b(r)\cos(n\theta)$ and $\Omega=w(r)\sin(n\theta)$. The level lines of $\Omega$ in the case $n=2$ are depicted on the right of Figure 3.
@@
If $n=1$, the homogeneous differential equation (A.9) with $b=0$ has a nontrivial solution $\varphi=rv_{*}$ which satisfies the boundary conditions. As a consequence, the inhomogeneous equation can be solved only if the right-hand side satisfies $\int_{0}^{\infty}b(r)r^{2}\,\mathrm{d}r=0$, and the solution is never unique. The solvability condition ensures that $F$ belongs to the subspace $\mathcal{Y}_{1}^{\prime}$ defined by